OpenStudy (anonymous):

the sum of two numbers is 50 and the diff is 24. what are the numbers?

OpenStudy (bgrg007):

37 and 13

OpenStudy (raden):

one of that number is (50+24)/2 = 74/2 = 37

OpenStudy (unklerhaukus):

A+B = 50 A = 50-B A-B =24 (50-B)-B = 24 50-2B = 24 50-24 = 2B

OpenStudy (anonymous):

so 26=2B so B=???? 13

hero (hero):

x + y = 50 x - y = 24 2x = 74 x = 74/2 x = 37 x = 13